OpenStudy (anonymous):

Find the other end point, Mid (4,-13) end (2,7)

OpenStudy (anonymous):



OpenStudy (anonymous):

Let A(x1, y1) be the required point, (4, -13)= (x,y) and (2,7) = (x2, y2) Now by Mid-point formula we have: \[\huge x= \frac{x_1+x_2}{2}, y= \frac{y_1+y_2}{2}\] Substituting the abovevalues we find: \[\huge 4= \frac{x_1+2}{2}, -13= \frac{y_1+7}{2}\] i.e 8 = x1+ 2 and -26 = y1+7 i.e. 8-2 =x1 and -26-7 =y1 i.e. 6= x1 and -33 = y1 i.e. x1= 6 and y1 = -33 hence A(6, -33) are the coordinates of the required other end point.

OpenStudy (anonymous):

@Alexjboy123